# Pilot Study on Generative AI and Critical Thinking in Higher Education Classrooms

### v1

@@ VI Acknowledgments
We would like to thank Annie Hui for her advice regarding the formulation and setup of this study. Emilia E We also want to thank the Stearns Center and the Computational and Data Sciences Department at GMU for their financial support.
@@ VII Author Contributions
Dr. Sharmin Abdullah is a corresponding author of this study. She contributed to the research by running CDS 130 classes during Summer Session C, guiding students and teaching assistants in conducting the necessary experiments and tasks for collecting data. She completed the required IRB training for this project and supported the development of the study through reviewing and editing both the manuscript and related funding proposals.

### v3

@@ VI Acknowledgments
We would like to thank Annie Hui for her advice regarding the formulation and setup of this study. Emilia Ermanoski in a special way, as her help with overseeing the TAs in the CDS 130 sections was instramental. We also want to thank the Stearns Center and the Computational and Data Sciences Department at GMU for their financial support.
@@ VII Author Contributions
Dr. Sharmin Abdullah contributed to the research by running CDS 130 classes during Summer Session C, guiding students and teaching assistants in conducting the necessary experiments and tasks for collecting data. She completed the required IRB training for this project and supported the development of the study through reviewing and editing both the manuscript and related funding proposals.